OpenStudy (anonymous):

What is the value of the expression |a+b| + |c| when a = -3, b = 7, and c = -15? A. -11 B. 5 C. 19 D. 25

OpenStudy (anonymous):

answer is C 19

OpenStudy (anonymous):

\[\left| a+b \right|+\left| c \right|\]

OpenStudy (anonymous):

\[\left| -3+7 \right|+\left| -15 \right|\]

OpenStudy (anonymous):

\[\left| 4 \right|+\left| 15 \right|\]